# Learning parameter dependence for Fourier-based option pricing with tensor trains

### v6

@@ Abstract
A long-standing issue in mathematical finance is the speed-up of option pricing, especially for multi-asset options. A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor trains to compress high-dimensional tensors. Another usage of the tensor train is to compress functions, including their parameter dependence. Here, we propose a pricing method, where, by a tensor train learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters. As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices. We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{5}$ and $10^{6}$ paths in terms of computational complexity while keeping comparable accuracy.
@@ I Introduction
Pricing options appropriately is needed for making a profit and managing the risk of loss in option trading. According to the theory in mathematical finance  [^1] As typical textbooks in this area, we refer to Refs. [1, 2]., the price of an option is given by the expectation of the discounted payoff in the contract with some stochastic model on the dynamics of underlying asset prices assumed. Except for limited cases with simple contract conditions and models, the analytical formula for the option price is not available and thus we need to resort to the numerical calculation. In the rapidly changing financial market, quick and accurate pricing is vital in option trading, but it is a challenging task, for which long-lasting research has been made. In particular, pricing multi-asset options, whose payoff depends on the prices of multiple underlying assets, is often demanding. Many pricing methodologies suffer from the so-called curse of dimensionality, which means the exponential increase of computational complexity with respect to the asset number, and the Monte Carlo method, which may evade the exponential complexity, has a slow convergence rate.
@@
In light of this, applications of quantum-inspired classical algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present  [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9]  [^3] Tensor train is a mathematically equivalent to matrix product state (MPS) [10]. learning to the Fourier transform (FT)-based option pricing method [11, 12], following the original proposal in Ref. [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^4] See [13, 14] as reviews. and recently, it has also been utilized in various fields such as machine learning [15, 16, 17], quantum field theory [18, 19, 20], and partial differential equations [7, 21, 22, 23, 24]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [25, 26, 19, 27] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
To address this issue, we focus on another usage of tensor trains that can embed parameter dependence [28, 29, 30, 31] to make FT-based option pricing more efficient. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We use these tensor trains to evaluate the integral including parameter dependence and perform fast option pricing in response to various parameter changes (refer to FIG. 1). To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ and $10^{6}$ paths by a factor of $O(10^{4})$ and $O(10^{5})$, respectively (see Fig. 2). We also confirm numerically that in the tested cases, the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths, and, although worse in some cases, is overall comparable to that with $10^{6}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
@@ II.1.1 Tensor cross interpolation
Tensor cross interpolation (TCI) is a technique to compress tensors corresponding to discretized multivariate functions with a low-rank TT representation. Here, we consider a tensor that, with grid points set in $\mathbb{R}^{d}$, has entries $F_{x_{1},\ldots,x_{d}}$ equal to $F(x_{1},\ldots,x_{d})$, the values of a function $F$ on the grid points [^5] Although we here denote the indexed of the tensor and the variables of the function by the same symbols $x_{1},\ldots,x_{d}$ for illustrative presentation, we assume that, in reality, the grid points in $\mathbb{R}^{d}$ is labeled by integers and the indexes of the tensor denotes the integers.. Leaving the detailed explanation to Refs. [25, 26, 19, 27], we describe its outline. It learns a TT using the values of the target function $F_{x_{1},x_{2}\ldots,x_{d}}$ at indexes $(x_{1},x_{2},\cdots,x_{d})$ adaptively sampled according to the specific rules. TCI actively inserts adaptively chosen interpolation points (pivots) from the sample points to learn the TT, which can be seen as a type of active learning. It gives the estimated values of the function at points across the entire domain although we use only the function values at a small number of sample points in learning. This is the very advantage of TCI and is particularly useful for compressing target tensors with a vast number of elements, contrary to singular value decomposition (SVD) requiring access to the full tensor. Note that TCI is a heuristic method, which means its effectiveness heavily depends on the internal algorithm to choose the pivots and the initial set of points selected randomly.
@@ III.1 Fourier transform-based option pricing
Here, the even natural number $N$ is the number of the grid points in one dimension. $\vec{z}_{\mathrm{gr},\vec{j}}$ is the grid point specified by the integer vector $\vec{j}$ as
$\vec{z}_{\mathrm{gr},\vec{j}}:=(\eta_{1}(\vec{p})j_{1},\ldots,\eta_{d}(\vec{p})j_{d}),$ (14)
where $\eta_{m}$ is the integration step size in the $m$-th direction. $\eta_{m}$ is a hyperparameter that must be appropriately determined according to the parameter $\vec{p}$ and thus written as $\eta_{m}(\vec{p})$ in (14). In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec. V for the detail. $\Delta_{\rm vol}:=\Pi_{m=1}^{d}\eta_{m}$ is the volume element.
@@ IV.1 Outline
Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set. Using TCI, we obtain the TTs to approximate $\phi$ and $\hat{v}$, incorporating the parameter dependence of these functions. The outline is illustrated in FIG. 1. Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$. Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$. Then, as illustrated in FIG. 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$. We reduce the bond dimensions of these TTs using SVD. We then contract adjacent core tensors pairwise and obtain two tensor train operators (TTOs), also known as matrix product operators (MPOs), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$. We contract these TTOs along the index vector $\vec{j}$ as follows:
$\tilde{V}^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (17)
Note that this resultant TTO has only the diagonal elements with respect to the local indices $\vec{k}$. That is, when $k_{1}\neq k^{\prime}_{1},\ldots,k_{d}\neq k^{\prime}_{d}$, we set $\tilde{V}^{k_{1},\ldots,k_{d}}_{k^{\prime}_{1},\ldots,k^{\prime}_{d}}=0$. With this operation, the resultant tensor can be expressed as TT $\tilde{V}_{k_{1},\ldots,k_{d}}$. Then we optimize the bond dimensions of $\tilde{V}_{k_{1},\ldots,k_{d}}$ using SVD. Having this TT, we get the option price for the specified value of $\vec{p}$ as illustrated in FIG. 1 (b). By fixing the index $\vec{k}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
@@
Here, we note that the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ for $\hat{v}$ and its TT approximation $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{i}$ corresponding to $\vec{p}$, which is now $\vec{\sigma}$ or $\vec{S}_{0}$, the parameter not affecting $\hat{v}$. This is due to $\eta_{j}$ depending on $\vec{p}$: because of the change of the grid points in the $\vec{z}$ space with respect to $\vec{p}$, the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ depends on the indexes $k_{i}$.
We also mention the ordering of the local indices of the TT. Two core tensors in the TT that correspond to $z_{j}$ and $p_{j}$ of the same asset are arranged next to each other, i.e., the order is ($z_{1}p_{1}z_{2}p_{2}\cdots z_{d}p_{d}$). In the demonstration in Sec. V, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing. On the other hand, if the core tensors on $\vec{z}$ and those on $\vec{p}$ are completely separated, i.e.,($z_{1}z_{2}\cdots z_{d}p_{1}p_{2}\cdots p_{d}$), we have found that the accuracy of the option pricing get worse since TCI fails to learn this tensor trains. However, the optimal arrangement of the local indices may vary, for example, depending on the correlation matrix: intuitively, the core tensors corresponding to highly correlated assets should be placed nearby. Although we do not discuss it in detail, this is an important topic for future research.
@@
Figure 1: Fast option pricing based on TNs proposed in this study. In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD. We then contract the tensors associated with $p_{m}$ and $z_{m}$, resulting in two TTOs. We then take the contraction of these two TTOs with respect to the index vector $\vec{j}$ and compress the bond dimension of the resulting TT using SVD. In (b), we use this TT to perform fast option pricing for a specified parameter $\vec{p}$.
@@ V.1 Details
With respect to $\vec{p}$, on which the TTs learn the dependence of the functions $\phi$ and $\hat{v}$, we take the two test cases: $\vec{p}$ is the volatilities $\vec{\sigma}$ or the present asset prices $\vec{S}_{0}$. In the proposed method, we need to set the range in the space of $\vec{p}$ and the grid points in it. For each volatility $\sigma_{m}$, we set the range to $\sigma_{m}\in[0.15,0.25]$, where the center $\sigma_{m}=0.2$ is a typical value of Nikkei Stock Average Volatility Index [34] and the range width ±$\pm 0.05$ covers the changes in this index on most days. For each $S_{m,0}$, we set the range to $S_{m,0}\in[90,120]$, which corresponds to the 20% variation of the asset price centered at 100. The lower bound is set to not 80 but 90 because the price of the option we take as an example is negligibly small for $S_{0}<90$. For both $\sigma_{m}$ and $S_{m,0}$, we set 100 equally spaced grid points in the range, and so the total number of the grid points in the space of $\vec{p}$ is $100^{d}$.
@@
$T$ $r$ $K$ $S_{0}$ $\alpha$ $\rho_{mm}$ $\rho_{mn(m\neq n)}$ $N$
$1$ $0.01$ $100$ $100$ $5/d$ $1$ $1/3$ $50$
@@
Integral step size
The spread of the characteristic function changes depending on the values of the parameters $\vec{p}$ and the number of assets. Accordingly, it is necessary to adjust the grid interval $\eta_{m}$ for integration in the wavenumber space.
For example, consider the case where $\vec{p}=\vec{\sigma}$. From Eq. (10), we see that for the current parameter values in TABLE 1, $\phi$ takes non-negligible value only when $\sigma_{m}z_{m}\lesssim 1$ for each $m$. It is necessary to ensure that the grid points cover the whole of such a region with a sufficiently fine interval. Currently, as $z_{m}$ ranges from $-N\eta$ to $+N\eta$ with step size $\eta_{m}$, $\eta_{m}$ and $\sigma_{m}$ must satisfy $\sigma_{m}\eta_{m}\ll 1$ and $\sigma_{m}\times N\eta_{m}\gg 1$. Therefore, in general, we need the adjustment that $\eta_{m}$ increases when $\sigma$ decreases and vice versa. Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{m}$ set to some constant, and thus we did so. In the cases with $d=10,11$, $\eta_{m}$ is adjusted as
$\eta_{m}=\frac{a}{\sigma_{m}}+b$ (19)
To set the constants $a$ and $b$, we ran TT-based option pricing for several values of $\sigma_{m}$ adjusting $\eta_{m}$, found the pairs $(\sigma_{m},\eta_{m})$ that yields the accurate price, and adjust $a$ and $b$ so that Eq. (19) approximately reproduces such $\eta_{m}$ from $\sigma_{m}$.
Also when $\vec{p}=\vec{S}_{0}$, we set $\eta_{m}$ to some constant in almost all the test cases. Only for $d=11$, we try the adjustment that
$\eta_{m}=\frac{\alpha}{S_{m,0}}+\beta$ (20)
with $\alpha$ and $\beta$ chosen according to the option pricing result for several $\eta_{m}$ and $\sigma_{m}$. In general, larger $S_{m,0}$, which means larger $x_{m,0}$, makes $\phi$ as a function of $z_{m}$ oscillate more rapidly, and thus we need finer grid points to resolve such an oscillation.
Note that when $\eta_{m}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$.
@@
We do not have the exact price of the multivariate min-call option since there is no known analytic formula for it. Instead, we regard the option price computed by the MC-based method with very many paths, concretely $5\times 10^{7}$ paths, as the true value. The error of the option price computed by the proposed method is evaluated by its deviation from the true value. We take the MC-based method with $10^{5}$ paths, which is a typical number in option pricing in practical business, as the comparison target against our method. We also run the MC method with $10^{6}$ paths, as the value when high accuracy is needed. As the error of the MC-based option price, we compute the half-width of its 95% confidence interval, namely $1.96\sigma/\sqrt{n_{\rm MC}}$, where $\sigma$ is the sample standard deviation of the discounted payoff in the MC run and the path number $n_{\rm MC}$ is $10^{5}$ or $10^{6}$. We assess the accuracy of our method by seeing if it falls within the statistical error range of the MC-based method.
An issue is that the number of possible combinations of the parameter $\vec{p}$ is $100^{d}$, and thus we cannot test all of them. Thus, we randomly select 100 combinations and perform option pricing for each of them. We compare the maximum among the errors of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
@@
TensorCrossInterpolation.jl [35] was used for learning tensor trains from functions appearing Fourier-based option pricing. ITensors.jl [36] was utilized for computing the contraction and inner product of the tensor trains. The Monte Carlo simulations were carried out using tf-quant-finance [37]. Parallelization was not employed in either case. GPUs were not utilized in any of the calculations. The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.

### v7

@@ Abstract
A long-standing issue in mathematical finance is the speed-up of option pricing, especially for multi-asset options. A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor trains to compress high-dimensional tensors. Another usage of the tensor train is to compress functions, including their parameter dependence. Here, we propose a pricing method, where, by a tensor train learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters. As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices. We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{6}$ paths in terms of computational complexity while keeping better accuracy.
@@ I Introduction
Pricing options appropriately is needed for making a profit and managing the risk of loss in option trading. According to the theory in mathematical finance  [^1] As typical textbooks in this area, we refer to Refs. [1, 2]., the price of an option is given by the expectation of the discounted payoff in the contract with some stochastic model on the dynamics of underlying asset prices assumed. Except for limited cases with simple contract conditions and models, the analytical formula for the option price is not available and thus we need to resort to the numerical calculation. In the rapidly changing financial market, quick and accurate pricing is vital in option trading, but it is a challenging task, for which long-lasting research has been made. In particular, pricing multi-asset options, whose payoff depends on the prices of multiple underlying assets, is often demanding. Many pricing methodologies suffer from the so-called curse of dimensionality, which means the exponential increase of computational complexity with respect to the asset number, and the Monte Carlo method, which may evade the exponential complexity, has a slow convergence rate. The quasi-Monte Carlo method has a better convergence rate in low-dimensional cases, but it can deteriorate in high-dimensional cases.
@@
In light of this, applications of quantum-inspired classical algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present  [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9]  [^3] Tensor train is a mathematically equivalent to matrix product state (MPS) [10]. learning to the Fourier transform (FT)-based option pricing method [11, 12], following the original proposal in Ref. [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^4] See [13, 14] as reviews. and recently, it has also been utilized in various fields such as machine learning [15, 16, 17], quantum field theory [18, 19, 20], and partial differential equations [7, 21, 22, 23, 24]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing  [^5] As another approach for accelerating FT-based multi-asset option pricing, see [25].. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [26, 27, 19, 28] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
To address this issue, we focus on another usage of tensor trains that can embed parameter dependence [29, 30, 31, 32] to make FT-based option pricing more efficient. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We use these tensor trains to evaluate the integral including parameter dependence and perform fast option pricing in response to various parameter changes (refer to FIG. 1). To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{6}$ paths by a factor of $O(10^{5})$ (see Fig. 2). We also confirm numerically that in the tested cases, the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{6}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the better accuracy.
@@ II.1.1 Tensor cross interpolation
Tensor cross interpolation (TCI) is a technique to compress tensors corresponding to discretized multivariate functions with a low-rank TT representation. Here, we consider a tensor that, with grid points set in $\mathbb{R}^{d}$, has entries $F_{x_{1},\ldots,x_{d}}$ equal to $F(x_{1},\ldots,x_{d})$, the values of a function $F$ on the grid points [^6] Although we here denote the indexes of the tensor and the variables of the function by the same symbols $x_{1},\ldots,x_{d}$ for illustrative presentation, we assume that, in reality, the grid points in $\mathbb{R}^{d}$ is labeled by integers and the indexes of the tensor denotes the integers.. Leaving the detailed explanation to Refs. [26, 27, 19, 28], we describe its outline. It learns a TT using the values of the target function $F_{x_{1},x_{2}\ldots,x_{d}}$ at indexes $(x_{1},x_{2},\cdots,x_{d})$ adaptively sampled according to the specific rules. TCI actively inserts adaptively chosen interpolation points (pivots) from the sample points to learn the TT, which can be seen as a type of active learning. It gives the estimated values of the function at points across the entire domain although we use only the function values at a small number of sample points in learning. This is the very advantage of TCI and is particularly useful for compressing target tensors with a vast number of elements, contrary to singular value decomposition (SVD) requiring access to the full tensor. Note that TCI is a heuristic method, which means its effectiveness heavily depends on the internal algorithm to choose the pivots and the initial set of points selected randomly.
@@ III.1 Fourier transform-based option pricing
Here, the even natural number $N$ is the number of the grid points in one dimension,$\vec{z}_{\mathrm{gr},\vec{j}}$ is the grid point specified by the integer vector $\vec{j}$ as
$\vec{z}_{\mathrm{gr},\vec{j}}:=(\eta j_{1},\ldots,\eta j_{d}),$ (14)
where $\eta$ is the integration step size in each direction, and $\Delta_{\rm vol}:=\eta^{d}$ is the volume element. $\eta$ is a hyperparameter that must be appropriately determined. In our demonstration, it is set to 0.4 or 0.3, which yields the accurate result (see Sec. V for details).
@@ IV.1 Outline
Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set. Using TCI, we obtain the TTs to approximate $\phi$ incorporating the parameter dependence of this function and $\hat{v}$. The outline is illustrated in FIG. 1. Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$. Considering $\phi$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$. Then, as illustrated in FIG. 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},\ldots,j_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ and $(\vec{z}_{\mathrm{gr},\vec{j}})$, respectively. We reduce the bond dimensions of these TTs using SVD. We then contract adjacent core tensors pairwise and obtain a tensor train operators (TTO), also known as matrix product operators (MPO), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$. We contract this TTO $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$. and TT $\tilde{v}_{j_{1},\ldots,j_{d}}$ along the index vector $\vec{j}$ as follows:
$\tilde{V}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}},$ (17)
Then we optimize the bond dimensions of $\tilde{V}_{k_{1},\ldots,k_{d}}$ using SVD. Thus, we get a new function $\tilde{V}_{k_{1},\ldots,k_{d}}$ whose input are parameter and output is option price. Having this TT, we get the option price for the specified value of $\vec{p}$ as illustrated in FIG. 1 (b). By fixing the index $\vec{k}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
@@
Here, we mention the ordering of the local indices of the TT. Two core tensors in the TT that correspond to $z_{j}$ and $p_{j}$ of the same asset are arranged next to each other, i.e., the order is ($z_{1}p_{1}z_{2}p_{2}\cdots z_{d}p_{d}$). In the demonstration in Sec. V, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing. On the other hand, if the core tensors on $\vec{z}$ and those on $\vec{p}$ are completely separated, i.e.,($z_{1}z_{2}\cdots z_{d}p_{1}p_{2}\cdots p_{d}$), we have found that the accuracy of the option pricing get worse since TCI fails to learn this tensor trains. However, the optimal arrangement of the local indices may vary, for example, depending on the correlation matrix: intuitively, the core tensors corresponding to highly correlated assets should be placed nearby. Although we do not discuss it in detail, this is an important topic for future research.
@@
Figure 1: Fast option pricing based on TNs proposed in this study. In (a), we learn TTs from parameter-dependent function of $\phi$ and $\hat{v}$ using TCI and reduce the bond dimension of these TTs using SVD. We then contract the tensors associated with $p_{m}$ and $z_{m}$ for TT of $\phi$, resulting in TTO. We then take the contraction of the TTO and TT with respect to the index vector $\vec{j}$ and compress the bond dimension of the resulting TT using SVD. In (b), we use this TT to perform fast option pricing for a specified parameter $\vec{p}$.
@@ V.1 Details
With respect to $\vec{p}$, on which the TTs learn the dependence of the functions $\phi$ and $\hat{v}$, we take the two test cases: $\vec{p}$ is the volatilities $\vec{\sigma}$ or the present asset prices $\vec{S}_{0}$. In the proposed method, we need to set the range in the space of $\vec{p}$ and the grid points in it. For each volatility $\sigma_{m}$, we set the range to $\sigma_{m}\in[0.15,0.25)$, where the center $\sigma_{m}=0.2$ is a typical value of Nikkei Stock Average Volatility Index [35] and the range width ±$\pm 0.05$ covers the changes in this index on most days. For each $S_{m,0}$, we set the range to $S_{m,0}\in[90,120)$, which corresponds to the 20% variation of the asset price centered at 100. The lower bound is set to not 80 but 90 because the price of the option we take as an example is negligibly small for $S_{0}<90$. For both $\sigma_{m}$ and $S_{m,0}$, we set 100 equally spaced grid points in the range, and so the total number of the grid points in the space of $\vec{p}$ is $100^{d}$.
@@
$T$ $r$ $K$ $S_{0}$ $\alpha$ $\rho_{mn},m\neq n$ $(N,\eta)$
$1$ $0.01$ $100$ $100$ $\frac{5}{d}$ $\frac{1}{3}$$\left\{\begin{array}[]{ll}(200,0.3)&;\ d=11,\vec{p}=\vec{S}_{0}\\ (100,0.4)&;\ \mathrm{otherwise}\end{array}\right.$
@@
We do not have the exact price of the multivariate min-call option since there is no known analytic formula for it. Instead, we regard the option price computed by the MC-based method with very many paths, concretely $5\times 10^{7}$ paths, as the true value. The error of the option price computed by the proposed method is evaluated by calculating the mean absolute error from the true value. As a comparison target for our method, we use an MC-based approach with $10^{6}$ paths. As the error of the MC-based option price, we compute the half-width of its 95% confidence interval, namely $1.96\sigma/\sqrt{n_{\rm MC}}$, where $\sigma$ is the sample standard deviation of the discounted payoff in the MC run and the path number $n_{\rm MC}$ is $10^{6}$. We assess the accuracy of our method by verifying if its error is below that of the MC-based method.
Here, an issue is that the number of possible combinations of the parameter $\vec{p}$ is $100^{d}$, and thus we cannot test all of them. Thus, we randomly select 100 combinations and perform option pricing for each of them. We compare the mean absolute error of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
@@
TensorCrossInterpolation.jl [36] was used for learning tensor trains from functions appearing Fourier-based option pricing. ITensors.jl [37] was utilized for computing the contraction and inner product of the tensor trains. The Monte Carlo simulations were carried out using tf-quant-finance [38]. Parallelization was not employed in either case. GPUs were not utilized in any of the calculations. The computations were performed on a 2023 MacBook Pro featuring a 12-core Apple M2 Max processor and 32GB of 400GB/s unified memory.